Two in-phase sources of waves are separated by a distance of 3.77 m. These sources produce identical waves that have a wavelength of 5.25 m. On the line between them, there are two places at which the same type of interference occurs. (a) Is it constructive or destructive interference? (b) and (c) Where are these places located (the smaller distance should be the answer to (b))?
Answer #1

a) distructive interference.

b) let r1 is the distance from source 1 and r2 is the distance from source 2 where distructive interefrence takes place.


minimum path diffrence for distructive intereference,

r1 - r2 = lamda/2

r1 - (3.77 - r1) = 5.25/2

2*r1 - 3.77 = 2.625

2*r1 = 2.625 + 3.77

r1 = (2.625 + 3.77)/2

= 3.1975 m

r2 = 3.77 - 3.1975

= 0.5725 m

so answer for part b) 0.5725 m

c) 3.1975 m